OpenStudy (help_people):

A polynomial has been factored, as shown below: f(x) = (x − 9)(x + 12)(x − 24) What are the zeros of the polynomial? −9, 9, −12, 12, −24, 24 −9, 12, −24 9, −12, 24 9, 12, 24

OpenStudy (help_people):

c? @welshfella @ganeshie8

OpenStudy (anonymous):

In order for f(x) to be zero, one of the factors must be zero (the only way for a product to be zero is if one or more of the factors is zero). So, to find the roots(zeros), set each factor equal to zero and solve for \(x\). \[x-9=0\]\[x+12=0\]\[x-24=0\]The solutions to these equations are the roots(zeros) of the function.

OpenStudy (help_people):

x-9=0 +9 x=9

OpenStudy (help_people):

could you make sure a few others are coreredct?

OpenStudy (help_people):

@Mehek14 @ospreytriple

OpenStudy (anonymous):

What did you get for the answer to this question?

OpenStudy (help_people):

i got c @ospreytriple @Mehek14

OpenStudy (help_people):

i know but osprey qas asking

OpenStudy (help_people):

and coudl you help me with a few more to maek sure they are ocrrect?

OpenStudy (anonymous):

Sorry. Not sure why the asker's answer doesn't appear in my thread.

OpenStudy (help_people):

thats ok :)

OpenStudy (help_people):

it happens all the time w/ me